OpenStudy (anonymous):

what number must you add in order to make x^2-4x=1 a complete square?

OpenStudy (anonymous):

I need helpe with the process.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

gblover, patiently go through this write-up and post ur doubts...we'll clear it. see basically the logic is to bring it to the form (x+b)^2=x^2+b2+2xb, so for getting this factorisation, see appropriately what to add. observe that the x term has co-efficient 2b and so if u hav a b^2 term u will form the complete square by the formula rite? so u add b^2

OpenStudy (anonymous):

here the 2b term is -4 and hence the b term is 2b=-4 and so b=-2. so u need to add (-2)^2 =4 get it?

OpenStudy (anonymous):

ok the process goes like this you have \[x^2+4x=1\] the coefficient of the x term is 4 1) half of 4 is 2 2) 2 squared is 4 so you add 4 to both sides and get \[x^2+4x+4=1+4\] \[x^2+4x+4=5\] and the left side is a perfect square, namely \[(x+2)^2=5\]

OpenStudy (anonymous):

sorry i wrote + when it should have been - but the process is the same

OpenStudy (anonymous):

so you subract 4 from both sides?